# Multi-Order Wavelet Derivative Transform for Deep Time Series Forecasting

### v1

@@ C.5 Hyperparameter Sensitivity Analysis on Additional Datasets
C.6 Selection of the Wavelet Basis Function
@@
Table 14: Forecasting performance of different Wavelet bases on four ETT datasets (top) and four other benchmarks (bottom).
@@
C.7 Ablation Study
The proposed WaveTS framework is centered on the Wavelet Derivative Transform, which is adept at capturing multi-scale, time-aware patterns. Specifically, the derivation in the frequency domain not only reflects stationary and robust frequency features but also reveals hierarchical characteristics that depict both low-frequency trends and high-frequency fluctuations concurrently. To validate the effectiveness of the derivation based on the Wavelet Transform, we conduct an ablation study, with the results summarized in Table 15. We design two forecasting models, namely FourierTS and WaveletTS, and conduct long-term time series forecasting experiments on four ETT and PEMS datasets. The architectures of these models for ablation studies are illustrated in Figure 4 where the Discrete Fourier Transform (DFT) and Discrete Wavelet Transform (DWT) are utilized as substitutes for the Wavelet Derivative Transform in the WaveTS framework. $\bm{\mathbf{H}_{t}}$ represents the frequency spectrum following the application of the DFT, while $\bm{\mathbf{LH}_{t}}$ and $\bm{\mathbf{LL}_{t}}$ detail the coefficient and approximation coefficient obtained after the implementation of the DWT. Instance Normalization and Frequency Refinement Units employed are consistent with those utilized in the WaveTS model. As summarized in Table 15, the WDT significantly enhances performance across various forecasting horizons, underscoring its efficacy in time series modeling.
Table 15: Ablation studies of the Wavelet Derivative Transform on four ETT and PEMS datasets in long-term forecasting. “I/O" indicates lookback window size/forecasting length. FourierTS and WaveletTS denote the substitution of the Wavelet Derivative Transform with the Fourier Transform and the Discrete Wavelet Transform based on our proposed WaveTS, respectively.
@@ C.7 Ablation Study
Figure 9: Architecture of the FourierTS and WaveletTS for ablation study where DFT and DWT replaces the WDT in WaveTS.
Appendix D Visualizations
D.1 Visualizations of the Forecasting Performance
(a) WaveTS
(b) FITS
(c) DeRiTS
(d) TexFilter
(e) iTransformer
(f) DLinear
Figure 10: Visualizations of the results on the ETTh1 dataset with lookback window size of 360 and forecasting length of 96.
(a) WaveTS
(b) FITS
(c) DeRiTS
(d) TexFilter
(e) iTransformer
(f) DLinear
Figure 11: Visualizations of the results on the Electricity dataset with lookback window size of 720 and forecasting length of 96.
D.2 Additional Visualizations of the Coefficients in the WDT
Figure 12: Additional visualizations of the Approximation Coefficients (LL) and Detail Coefficients (LH) in the WDT on four different datasets. We select the “OT” variate with the input length is 720 for each dataset.
Derivative of the sequence is essential for identifying more semantically rich and robust representations. As illustrated in Figure 12, LL reveals macro temporal patterns, providing insight into the overall trend, while LH captures detailed fluctuations and complex irregular micro patterns at various scales. Prior research on frequency derivative learning indicates that derivation yields more stationary frequency patterns, as depicted in Figure 5 in [7]. However, the conversion of the original sequence into stationary representations may neglect to capture the overall trend, the most fundamental and informative temporal pattern indicating nonstationarity, thereby resulting in ineffective modeling. As demonstrated in Figure 12, WaveTS effectively addresses this problem by decomposing the original signal into LL and LH coefficients, wherein the former encapsulates trend and periodicity, while the latter captures more stationary characteristics. Consequently, WaveTS is capable of modeling hierarchical and robust frequency features simultaneously, which significantly enhances forecast performance.
D.3 Additional Visualizations of the Abrupt Changes Amplification
Figure 13: Additional visual evidence that the Wavelet Derived Coefficients (red) produced by the proposed WDT amplify abrupt regime shifts more clearly than the standard Wavelet Coefficients (blue). For four representative datasets, the black dashed boxes mark segments where the original series (black) undergoes sudden trend reversals or variance bursts. In every case, sharp spikes or sign reversals appear in the WDT coefficients, whereas the plain WT responds more mildly, confirming that WDT magnifies rate-of-change cues.
Across all four datasets, the highlighted windows show that the magnitude of the wavelet derived coefficients surges precisely when the original series experiences a swift trend break and volatility jump. This behavior is consistent with the analytical role of the derivative operation as a local high-pass filter: it suppresses slow-moving components while accentuating rapid derivatives. By contrast, the standard WT coefficients remain comparatively smooth, indicating limited sensitivity to such localized shocks. The repetition of this pattern on time series collected from various domain demonstrates that the WDT generalizes beyond a single domain, providing a robust mechanism for exposing change points that would otherwise be under-represented in frequency representations.
D.4 Additional Visualizations of the Time-Frequency Scalogram
@@
Appendix E Limitations
While WaveTS achieves state-of-the-art accuracy and efficiency on a diverse collection of regularly-sampled benchmarks, two avenues for improvement remain. First, our empirical study is confined to fixed-interval datasets; extending the model to irregularly-sampled or event-driven series would clarify its robustness in real-world monitoring scenarios. Second, the current implementation selects the derivative order and wavelet scale via grid search; developing an adaptive mechanism—e.g., a learnable branch-selection policy—could further reduce tuning overhead and improve scalability to longer horizons without enlarging the architectural footprint.

### v2

@@ D.5 Hyperparameter Sensitivity Analysis on Additional Datasets
D.6 Fair and Reproducible Hyperparameter Selection
Lookback Window Length $L$.
We treat $L$ as a tunable hyperparameter because architectures require different effective receptive fields to perform optimally. Enforcing a single $L$ (e.g., 96) conflates equality with fairness and disadvantages models (e.g., DLinear (Zeng et al., 2023)) that benefit from longer histories, while many Transformer variants gain little from longer inputs. For WaveTS, a larger $L$ improves time–frequency fidelity and remains affordable; thus we tune $L$ on the validation set. For baselines, we strictly follow lookbacks from their papers/official code (defaults when unspecified). Evaluating each method under its own validated configuration is widely accepted and improves fairness and reproducibility (Zeng et al., 2023; Yang et al., 2024; Xu et al., 2024).
WDT Order $O$ and Scale $S$.
WaveTS uses the WDT derivative order $O$ (number of branches $N$ in Eq. 3) and scale $S$ (decomposition levels $K$ in Eq. 3). For each dataset and horizon, we pick $(O,S)$ via a tiny, validation-driven grid (optionally with Optuna), typically $O,S\in\{1,\ldots,5\}$. There is no closed-form rule from data traits to optimal $(O,S)$, so the search is principled, lightweight, and reproducible. Sensitivity analyses in Figure 3 and Figure 8 show performance is flat around the optimum: periodic/stable datasets (e.g., Traffic, Electricity) often prefer smaller $O,S$; more non-stationary, multi-scale datasets benefit from larger $O,S$; longer horizons favor a larger $S$.
D.7 Additional analysis on the order and the scale
Contribution of each branch. As described in Section 3.2.2, after obtaining the multi-branch time-domain representations $\{\mathbf{Z}_{n}\}_{n=1}^{N}$ in Eq. 5, we concatenate them along the time dimension and apply a linear projection to obtain the final reconstruction and prediction as in Eq. 6. To make the role of each derivative-order branch explicit, we provide a quantitative analysis of its contribution to the final forecast. Because the final aggregation is strictly linear, the projection weight in Eq. 6 can be written as the block matrix $\mathbf{W}=[\mathbf{W}^{(1)}\;\mathbf{W}^{(2)}\;\cdots\;\mathbf{W}^{(N)}]$, where each block $\mathbf{W}^{(n)}$ multiplies the corresponding branch output $\mathbf{Z}_{n}$. This yields the exact additive decomposition $\mathbf{Y}_{t}=\sum_{n=1}^{N}\mathbf{W}^{(n)}\mathbf{Z}_{n}+\mathbf{b}$, and we define the per-branch contribution as $\mathbf{Y}_{t}^{(n)}=\mathbf{W}^{(n)}\mathbf{Z}_{n}$, so that $\mathbf{Y}_{t}=\sum_{n=1}^{N}\mathbf{Y}_{t}^{(n)}+\mathbf{b}$. Since no nonlinearity is applied after concatenation, there are no cross terms and the superposition principle holds. We quantify each branch’s impact using the $\ell_{2}$ energy share $\rho_{n}^{(2)}=\frac{\lVert\mathbf{Y}_{t}^{(n)}\rVert_{2}}{\sum_{m=1}^{N}\lVert\mathbf{Y}_{t}^{(m)}\rVert_{2}}$, computed on the test set. This incurs no retraining and only a single forward pass, together with a closed-form split of $\mathbf{W}$ into its $N$ column blocks $\{\mathbf{W}^{(n)}\}$. Tables 14 and 15 report $\rho_{n}^{(2)}$ for WaveTS with four branches $(O=4)$ on the Traffic and ETTm2 datasets across horizons $\{96,192,336,720\}$. Lower-order branches dominate on strongly periodic data (Traffic), whereas higher-order branches contribute more on more nonstationary data (ETTm2), which is consistent with our hyperparameter selection rationale. Boldface marks the largest contribution.
Table 14: $\ell_{2}$ energy share on Traffic (larger indicates a greater contribution).
Traffic_96 Traffic_192 Traffic_336 Traffic_720
Branch 1 $(O=1)$ 0.301543 0.299975 0.316984 0.264803
Branch 2 $(O=2)$ 0.281398 0.250889 0.254451 0.276797
Branch 3 $(O=3)$ 0.217832 0.244350 0.242937 0.219048
Branch 4 $(O=4)$ 0.199227 0.204786 0.185628 0.249352
Table 15: $\ell_{2}$ energy share on ETTm2 (larger indicates a greater contribution).
ETTm2_96 ETTm2_192 ETTm2_336 ETTm2 _720
Branch 1 $(O=1)$ 0.242765 0.244528 0.237057 0.250936
Branch 2 $(O=2)$ 0.242479 0.228037 0.254821 0.222338
Branch 3 $(O=3)$ 0.251197 0.264088 0.230385 0.264954
Branch 4 $(O=4)$ 0.263559 0.263347 0.277738 0.261772
Which orders and scales suit which datasets? Complementing the contribution analysis above, we examine which derivative order $O$ and wavelet scale $S$ are suitable for different datasets and horizons. In our setting, $O$ equals the WDT derivative order and $S$ is the number of wavelet decomposition levels. For each dataset and each forecasting horizon, we select $(O,S)$ on the validation set via a small grid (optionally accelerated by Optuna), typically $O,S\in\{1,\dots,5\}$. We adopt this protocol because there is no closed form mapping from dataset characteristics such as periodicity and nonstationarity, together with horizon length, to the optimal $(O,S)$. The search space is tiny, so the procedure is negligible in cost and reproducible; moreover, our sensitivity analyses show performance is relatively flat near the optimum, indicating the search chooses among near equivalent settings rather than overfitting. Empirically, the validated choices in Tables 16 and 17 follow a consistent pattern: strongly periodic datasets such as Traffic and Electricity tend to prefer smaller $O$ and $S$, while datasets with more pronounced nonstationarity or multi scale fluctuations such as ETTm1 and ETTm2 often benefit from larger $O$ and $S$. Longer horizons usually favor a slightly larger $S$ to stabilize low frequency forecasting.
Table 16: Traffic and Electricity.
Traffic Electricity
Horizon $O$ $S$ Horizon $O$ $S$
96 2 3 96 2 2
192 3 2 192 3 2
336 3 4 336 4 3
720 2 4 720 3 5
Table 17: ETTm1 and ETTm2.
ETTm1 ETTm2
Horizon $O$ $S$ Horizon $O$ $S$
96 3 2 96 2 2
192 3 3 192 4 3
336 4 5 336 4 4
720 5 5 720 4 5
D.8 Selection of the Wavelet Basis Function
@@
Table 18: Forecasting performance of different Wavelet bases on eight datasets .
@@
D.9 Ablation study
The WaveTS framework is centered on the Wavelet Derivative Transform (WDT), which captures multi scale and time aware structure by separating low frequency trends and high frequency fluctuations while preserving exact invertibility. To isolate the effect of the front end transform, we design two ablated models, FourierTS and WaveletTS, by substituting the WDT in WaveTS with the Discrete Fourier Transform (DFT) and the Discrete Wavelet Transform (DWT), respectively, while keeping all other components and training settings identical (normalization, Frequency Refinement Units, linear projection, and losses). The ablation architectures are illustrated in Figure 9. Results on four ETT datasets and four PEMS datasets are summarized in Table 19, showing that WDT consistently improves long horizon forecasting across metrics and horizons.
[…]
Table 19: Ablation of the transform front end on four ETT datasets in long horizon forecasting. “I/O” indicates lookback window size and forecasting length. FourierTS and WaveletTS replace WDT with DFT and DWT in WaveTS, respectively.
@@ D.9 Ablation study
Figure 9: FourierTS and WaveletTS for the ablation, where DFT and DWT replace WDT in WaveTS.
Across all datasets and horizons in Table 19, WaveTS surpasses both FourierTS and WaveletTS. The gains are most pronounced at longer horizons, where derivative aware multi order branching provides richer low frequency stabilization and high frequency discrimination than a global Fourier spectrum or a derivative free wavelet stack. These findings support the design choice of coupling band localization with derivative emphasis in WaveTS.
Appendix E Visualizations
E.1 Visualizations of the Forecasting Performance
(a) Visualizations of the results on the ETTh1 dataset with lookback window size of 360 and forecasting length of 96.
(b) Visualizations of the results on the Electricity dataset with lookback window size of 720 and forecasting length of 96.
E.2 Additional Visualizations of the Coefficients in the WDT
Figure 15: Additional visualizations of the Approximation Coefficients ($L$) and Detail Coefficients (LH) in the WDT on four different datasets. We select the “OT” variate with the input length is 720 for each dataset.
Derivative of the sequence is essential for identifying more semantically rich and robust representations. As illustrated in Figure 15, LL reveals macro temporal patterns, providing insight into the overall trend, while LH captures detailed fluctuations and complex irregular micro patterns at various scales. Prior research on frequency derivative learning indicates that derivation yields more stationary frequency patterns, as depicted in Figure 5 in (Fan et al., 2014). However, the conversion of the original sequence into stationary representations may neglect to capture the overall trend, the most fundamental and informative temporal pattern indicating nonstationarity, thereby resulting in ineffective modeling. As demonstrated in Figure 15, WaveTS effectively addresses this problem by decomposing the original signal into LL and LH coefficients, wherein the former encapsulates trend and periodicity, while the latter captures more stationary characteristics. Consequently, WaveTS is capable of modeling hierarchical and robust frequency features simultaneously, which significantly enhances forecast performance.
E.3 Additional Visualizations of the Time-Frequency Scalogram
@@
Appendix F Limitations
While WaveTS achieves state-of-the-art accuracy and efficiency on a diverse collection of regularly-sampled benchmarks, two avenues for improvement remain. First, our empirical study is confined to fixed-interval datasets; extending the model to irregularly-sampled or event-driven series would clarify its robustness in real-world monitoring scenarios.
Appendix G The Use of Large Language Models (LLMs)
We employ LLMs solely to correct grammatical errors and to ensure the writing is fluent, accurate, and coherent. The authors take full responsibility for all content generated with LLM assistance.